IEPM

Status Report for SLAC/DOE IEPM Project

For February 2002

SLAC Home Page
Work completed or in progress since the last report.

Work Related to Analysis and Development

PingER

Warren worked with the team at FNAL to implement some new metrics on the archive graphing tool and added the IEPM-BW data to the PingER reporting tool. He also worked with the remaining monitoring sites still needing to upgrade to timeping 1.1.7 and new beacons list.

Warren and Les reviewed the 'continents' associated with the nodes monitored by PingER and reassigned many to better describe the region.

IEPM-BW / Grid-Monitoring

Les worked with LBNL, NASA/GSFC and the University of Wisconsin to resolve problems with IEPM-BW testing. Les worked with BNL to get IEPM-BW through their firewall.

Les worked with Margaret Murray of CAIDA and Brian Tierney of LBNL to understand and validate IEPM-BW results. Les also worked with Rick Wolski of NWS concerning making predictions, and worked with Connie and Manish to implement predictions into the code.

IPEX

A vulnerability was found in the SSHD in Cyclades Box used in the IPEX data gathering machine. Warren investigated the possibility of using a SLAC taylored machine to act as the IPEX machine. Security vetoed our plans to plug either into the network until a formal request was made and approved by the SLAC CCC.

Warren and Les had a telephone conference with some of the developers do discuss the approach to security/privacy.

Globus

Adil installed Globus on Antonia. Warren made some measurements with GridFTP.

Warren is working with UCL to install DOE-SG certificates to extend testing to Europe.

Visualization

Warren looked at using the UTH package to plot IEPM-BW and Web100 data.

IPMP

Waiting for security to reconfigure the ACLs and allow IPMP traffic between the SLAC AMP box and other hosts involved in the project.

VoIP

Warren worked with David Howard, the telephone syustem manager, to set up an answering machine for testing from Brazil.

Warren talked with Campus about our VoIP setup.

Missing Data

Transpac (now at Starlight) returned to service.

The Australia monitoring site is a no-hoper.

Worked with FNAL to fill gaps in their database.

IPv6

Warren received an enquiry about support for the IPv6 monitoring and reported this is no longer maintained due to lack of manpower. He also briefed Bob Fink on what is requried to run it and Bob will look for assistance from the Internet2 IPv6 WG.

Misc

The usual paperwork, installations and administration.

Work related to Publicity and Information

Paper for Comp Comm

Warren, Les and Connie worked on a paper and submitted it to the editors for a special edition of Computer Communications.

Internet2

Les spent 2 days in Tempe AZ, at the Internet 2 End to End performance initiative and the I2 HENP network working group.

PingER Report

Les wrote the PingER half yearly report.

INCITE Report

Les wrote a report on SLAC's contribution to the INCITE proposal.

GGF

Les presented talks on QOS ( PPT or PDF), network measurement ( PPT or PDF ), and IEPM-PPDG ( PPT or PDF) at GGF.

Created March 7, 2002.
Comments to iepm-l@slac.stanford.edu